You are here: Foswiki>Commissioning Web>Com20130903 (05 Sep 2013, TarasGolota)Edit Attach

UT 20130903 Technical Observing - LUCI1 and AGW1 Alignment at LFBG

Observer: JHill, DThompson (Tucson) DMiller (LBTO)
Telescope Operator: GBechetti (LBTO)
Software Support: DCox (Tucson)
AO Support: DMiller (LBTO)
Telescope Support: JUrban (LBTO)
Instrument Support: ESolheid (LBTO)

Summary

Dave took data for LUCI1 rotator center and zeropoint, and he is analyzing it overnight (see results below). If weather and seeing are good, we will start Tuesday evening with LUCI1 transformation data.

We took field aberration data for AGW1@LFBG and confirmed that binodal aberration is small (at both LFBG and LDG see results below).

We did the AGw2 wake-up activities and did the initial alignment. The alignment had to be a multi-step bootstrap process as M2 and M3 and AGW2 had all been dismounted.

Preparations

Details

Dave T. working on LUCI1 pupil alignment in twilight.

Open

02:31 Open to mostly clear, but not photometric sky. T=14.5 degC, D=5 degC

LUCI1 Rotator Zeropoint and Rotator Center

Dave T. is doing this in a separate log........

03:45 Dave T. is done with his data collection

DThompson wrote:
        From the 6 images (3 fields, 2 exposures each) I get a mean residual rotation 
(ignoring the usual significant digits rules) of 0.006 +/- 0.030 degrees,
so there is no need to change the LUCI1 LEFTZEROPOINT from the current value.

        For the rotator center in the N3.75 camera, I find it is at 1017.17  1040.66 with 
about 1-2 pixel rms.  This is comparable to where it usually lands (offset about 2 arcsec 
from the center of the detector).  Don't believe the info in the headers…that is the rotcen 
for the N1.8 camera.  You can only have one rotcen in the LUCI1 software.

03:50 Doug M is testing his remote collection of LUCI images (getting ready for transformation data).

04:00 Seeing degrades from 0.8 to 2 arcsec, as we go through some clouds. That puts our plans for transformation data on hold.

Doug is having lots of problems trying to display the LUCI desktop on obs3 as LBTO - we think it is related to problems with the monitor configuration. See Issue 4783.

Practice Transformation Data on StoneO

/home/lbto/data/20130903/transform_test/

Doug is collecting practice transformation data on the StoneO field. This is to make sure everything is working. He is also making some improvements to the scripts such as trapping errors returned by the request for a LUCI image.

05:25 The seeing is as good as 0.7 arcsec on the guider, but variable seeing and intermittent clouds make it difficult to get any consistent data.

Field Data on BS9177 using AGw1@LFBG

05:36 ACTIVE preset to BS9177 - seeing on guider is ~0.8 arcsec FWHM and variable - EL~57

/home/lbto/data/20130903/field_agw1/

05:43 field_collect

The seeing varied from 0.6 to 1.1 arcsec during this data set.

06:29 Seeing puffs up above 2 arcsec at image 000314 then back to 1.1 arcsec for the next image.

06:41 Seeing blows up to 2 arcsec again

06:44 Lost the star to clouds.......image 000326 is no good.

Getting this message a couple times per hour on the GCSGUI: "GCS is guiding but PCS is not accepting g"

see IT #4781 https://lbt.issuetrak.com/Issue_View.asp?IssueNbr=4781

06:55 Lost star due to clouds --- end of this data.

GStar wfsc #
0 275
11 276-278
0 279
16 280-282
0 283
13 284-286
0 293-289
17 294-296
0 297
21 298-300
0 301-302
6 303-305
0 306-307
3 308-310
0 311
4 312-314
0 315
12 316-318
0 319
8 320-322
0 323
18 324-326
0 327-328
11 329-331
0 332-333

07:21 Field data Analysis DLM
  • Could be some binodal but not really above the noise
  • First part of data set was OK (0.8" seeing) but latter data poor (~1.0" seeing)
  • No M2 position changes made

 => New Field Data Results
             Measured bi-nodal astigmatism
              From Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  -0.5", Delta IE =   0.5"
         Tilt =   1.9", Delta CA =  -1.9"
  
                M2 Mis-alignment 
             M2 Arcseconds and Millimeters
         -------------------------------------------
          Delta RX =    -5.", Delta Y = -0.025 mm
          Delta RY =   -19.", Delta X = -0.098 mm
  • LFBG Field Astigmatism:
    lfbg_field.jpg

Switch to AGW2@RFBG for aliveness testing

06:58 to zenith to reconfigure swing arms

07:25 Found star M50036 at EL=51.7 at IE=-104 CA=-177

OK, everything is alive.

Realignment after M2 has been dismounted.

See the procedure in BGTechProcedures

Now we have to bootstrap the alignment when AGW2 and M3 and M2 have been dismounted.

ptreset to the default IE=-43 CA=-124

Adjusting M2 with coma free pointing -- hit limits on M1 and M2 ?? range-balancing?

Can't move M2 coma-free tilt in the negative direction that we need to go, because we hit the RY=899 limit.

Adjust the coma first X=-2.8 Y=-4.65 RX=485 RY=-291 so that buys a bit of RY headroom.

CA can't go any farther than -147 with Tilt limited to -30.

Tip needs -61 to get IE back to -43.

Doug points out that the M3 has been remounted also, so this initial procedure for adjusting collimation after M2 has been dismounted is not valid. Next we work our way through a more bootstrap alignment procedure.....

Tip=-61 Tilt=-30 on M2 gives X=1.53 Y=3.11 RX=-599 RY=294

Doug and Geno trying active optics while John on Telecon with Italy. The active correction pushed M2 into the RY limit again.

09:04 Apply to M2 X=-2 RY=-200 Global offsets to get off the RY limit.

09:06 Apply John's coma correction from above to the M2 global offsets. X=-4.8 Y=-4.65 RX=485 RY=491

We do not have any repointing applied to the secondary at this point, so we are using the corrected IE/CA values.

Updating the on-axis pupil position on the wfs

09:09 ACTIVE Preset

First we need to update the on-axis pupil position on the wfs, and we need to adjust the guiding hotspot.

09:22 New on-axis pupil center is 359.6, 258.4 ---> Doug applied this to WFSCam_hotspot_x in LUCI_IRTC_R.cfg (takes effect without restarting GCSR)

09:29 Clear Primary figure corrections, as we had some high order stuff stuck in there. That makes the wavefront look substantially

09:30 Doug is lowering the gain coefficients for Z9, Z10, Z11 to match those in use on the left side.

Then we ran into the M2 -Y limit (really the 10 mm radius)

Problems with Active Optics running away when hotspot is incorrect.

Adjusting the guiding hotspot

09:54 Adjusting guiding hotspot to get light into the wfs ---> 280, 230 pixels unbinned

Measuring Pupil Wobble from DX M3

10:13 Have nice 0.8 arcsec image now. Begin measuring pupil wobble.

Total Collimation M2 is X=0.052 Y=-6.875 Z=0.744 RX=391 RY=536 at EL=58.8

but Doug's pupil_collect routine had the side of the preset and offset hardwired to left.......he fixed that in a couple minutes

0.00000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000002.fits
60.0000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000003.fits
120.000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000004.fits
180.000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000005.fits
240.000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000006.fits
300.000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000007.fits

First measurement was at Tip=250 Tilt=-150 Absolute -208 +529 -25.831

Wispy clouds moving over.....

10:33 Move DX M3 Global Offset Tip from +200 to +500 in steps of 50 allowing the guider to follow

collimate_once 008

/home/aoeng/data/20130903/pupil_wobble/

10:40 Move DX M3 Global Offset Tip from +500 to -300 arcsec. Tilt remains -150.

collimate_once 009

collimate_once 010

10:47 move DX M3 Global Offset Tilt from -150 to +50 arcsec. Tip remains -300.

collimate_once 011

10:50 move DX M3 Global Offset Tilt from +50 to +150 arcsec. Tip remains -300.

collimate_once 012

10:53 move DX M3 Global Offset Tilt from +150 to +250 arcsec. Tip remains -300.

collimate_once 013

10:56 Move DX M3 Global Offsets Tip from -300 to -200 arcsec. Tilt remains +250 arcsec
Absolute coordinates: Tip=-608 Tilt=+930 RZ=-25.831

JMH: These were archived in DXTMIRTCFCollimation.dat the following day.

(also adjusting focus and astigmatism by hand)

collimate_once 014 looks good

pupil_collect

10:58 Why did the guider image jump at the rotation offset of zero?

11:03 Lost star during the offset. Why?

11:05 ACQUIRE preset to M5 star IE=-114 CA=-139 (but that included some manual pointing offsets)

11:08 ACTIVE preset

Alignment of remounted M2

11:11 ACQUIRE preset to pointing star

ptreset 1

Clear global offsets; Clear manual pointing; Clear Active Optics XYZ

M2 Tip -80

Move CA to -139

Manually apply -N000 Z7 and +M000 Z8

M2 Active Optics X=-2.795 Y=-4.658 RX=485 RY=-291

ptreset back to default IE=-43 CA=-124

M2 Tip=-84 Tilt=-24 ---> Manual Pointing X=1.226 Y=+4.290 RX=-825 RY=236

Success! (It is within range.) Need to sum Active Optics and Manual Pointing for the new Global Offsets.
New Global Offsets for DX M2 collimation are: X=-1.569 Y=-0.368 Z=0 RX=-340 RY=-55

JMH- The following day: Those Global Offsets were installed in DXSMIRTCFCollimation.dat and all the other DXSM* files.

Beware that this right side collimation LUT could have a large binodal term with elevation that we need to remove.

Refinement of the Pupil Wobble adjustment

11:39 ACTIVE preset to BS -- seeing is down to 0.5 arcsec on guider

stopWFSing

11:45 pupil_collect

0.00000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000020.fits
60.0000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000021.fits
120.000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000022.fits
180.000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000023.fits
240.000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000024.fits
300.000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000025.fits

Results of above are peculiar, so we will collimate and try again. See the red curve in the plot below.

12:00 ACTIVE preset to clean up the pupil a bit.

12:03 pupil_collect

0.00000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000026.fits
60.0000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000027.fits
120.000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000028.fits
180.000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000029.fits
240.000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000030.fits
300.000 /home/aoeng/data/20130903/pupil_wobble/right_wfsc000031.fits

Results of above are also peculiar. The pupil locations are extended in the X-direction, but not Y.

  • Pupil Wobble at RFBG:
    pupil_wobble.jpg

Close

12:20 close for impending dawn

-- JohnHill - 03 Sep 2013
Topic attachments
I Attachment Action Size Date Who Comment
lfbg_field.jpgjpg lfbg_field.jpg manage 28 K 03 Sep 2013 - 07:20 DougMiller LFBG Field Astigmatism
pupil_wobble.jpgjpg pupil_wobble.jpg manage 181 K 03 Sep 2013 - 12:25 DougMiller Pupil Wobble at RFBG
Topic revision: r11 - 05 Sep 2013, TarasGolota
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback